OpenStudy (anonymous):

which of the following is equivalent to the expression: sin^2 theta(csc^2 theta-1)???

OpenStudy (anonymous):

A. sin^2 theta B. cos^2 theta C. csc^2 theta D. sec^2 theta

OpenStudy (anonymous):

Firstly, you know that: \[\csc\theta=\frac{1}{\sin\theta}\] What would: \[\csc^2\theta=?\]

OpenStudy (anonymous):

Remember: \[\csc^2\theta=(\csc\theta)^2\] So what would \[\csc^2\theta=?\] in terms of sintheta?

OpenStudy (anonymous):

I honestly don't know/:

OpenStudy (anonymous):

would it be hyp/opp?

OpenStudy (anonymous):

\[\csc\theta=\frac{1}{\sin\theta}\] Right? And if: \[\csc^2\theta=(\csc\theta)^2\] Then it also equals: \[=(\frac{1}{\sin\theta})^2\] Now get rid of the exponent and tell me what it equals?

OpenStudy (anonymous):

(1/sin theta)^2 ?

OpenStudy (anonymous):

That means: \[\frac{1}{\sin\theta}\times \frac{1}{\sin\theta}=?\]

OpenStudy (anonymous):

@pauli